A long-distance phone call costs $0.96 for the first minute and $0.72 for each additional minute or portion thereof. Write an inequality reprssenting the number of minutes a person could talk without exceeding $6.
Answer by ptaylor(2198)

x=number of minutes a person can talk without exceeding $6.00
$0.72(x-1)+$0.96 =cost for x minutes and that has to be less than or equal to $6.00. So, our inequality is:
$0.72(x-1)+$0.96=<$6.00 clear parens
$0.72x-$0.72+$0.96=<$6.00

$0.72x+$0.24=<$6.00 subtract $0.24 from both sides
$0.72x=<$6.00-$0.24
$0.72x=<$5.76 divide both sides by $0.72
x=<8 minutes-----------------------------answer

ck
8 minutes costs $0.96 + 7($0.72)=
8 minutes cost $0.96+$5.04=$6.00



Hope this helps-----ptaylor